OpenStudy (anonymous):

How can trigonometric ratios help a pilot and the ground crew maintain safety when flying?

OpenStudy (anonymous):

MEDALS MEDALS

OpenStudy (anonymous):

Someone help pleasee!

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

consider a mountain 3km in the direction of the airplane and height of 5km... with this values we can find the angle of slope for don't crash with the mountain... for example \[\tan \theta=\frac{ 5 }{ 3 }\] or \[\theta = \tan^{-1} \frac{ 5 }{ 3 }\]